OpenStudy (anonymous):

find limit as x approaches pi/2 from negative side. f(x)=(tanx)^[(pi/2)-x]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/2^{-}} tanx ^{(\pi/2)-x}\], answer should be "1" but i need to know the step by step process

OpenStudy (anonymous):

f(x) = tanx ^ [(pi/2) - x] f(x) = e^ [(pi/2 - x) x ln tanx ] when x approaches pi/2, then f(x) = e^0 = 1

OpenStudy (anonymous):

thank you jyouie :)

OpenStudy (anonymous):

Nope, what is your age now? This is something more advanced....

OpenStudy (anonymous):

actually i'm 20 and in college

OpenStudy (anonymous):

so are you saying thats not the answer? aha

OpenStudy (anonymous):

...I'm not quite sure, 'cause me too, I'm still studying at secondary school (something like high school in us)....18 ^^

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

well it seems like you know a lot more than I do, you haven't even taken the advance course yet. i give you kudos.

OpenStudy (anonymous):

you know how to use l'hopital's rule?